# Intrinsic superconducting diode effect and nonreciprocal superconductivity in rhombohedral graphene multilayers

### v1

@@ Abstract
One experimental manifestation of the asymmetric quasiparticle dispersion is that a transparent normal-metal–superconductor junction can exhibit a current-voltage ($I$-$V$) characteristic that is not invariant under $V\rightarrow-V$ [53]. Our results suggest the following probe of the nonreciprocal superconductivity in RTLG: connect two normal leads to the RTLG superconductor, one along $\theta=0^{\circ}$ ($\Gamma$-$K$ direction) and one along $\theta=90^{\circ}$ ($\Gamma$-$M$ direction). Our results imply that the lead along $\theta=0^{\circ}$ would show an asymmytric $I$-$V$ curve, while the $I$-$V$ curve along $\theta=90^{\circ}$ would remain symmetric.
@@
Intrinsic SDE. We are now in the position to evaluate the SDE for RTLG. For this purpose, we self-consistently solve for the superconducting order parameter for a broader range of Cooper-pair momenta, $\boldsymbol{q}$, and use the obtained $\Delta_{\boldsymbol{k},\boldsymbol{q}}$ to evaluate the free energy. The results of the free energy are shown in Fig. 3(a). The current in the ground state that minimizes $F(\mathbf{q})$ is always zero. However, applying a finite current-bias, produces a state with a finite Cooper-pair momentum different from the ground state value, $\boldsymbol{q}\neq\boldsymbol{q}_{0}$. The current in this state is $I=(2e/\hbar)|\partial_{\boldsymbol{q}}F(\boldsymbol{q})|$[13].
@@
As a next step, we compute the supercurrent versus $\boldsymbol{q}$ are shown in Fig. 3(b). Upon fixing a direction, $\theta$, so that $\boldsymbol{q}=(q\cos\theta,q\sin\theta)$, this allows us to determine the critical current in the forward direction, $I_{c,+}\equiv\text{max}_{q>0}\{I(q,\theta)\}$ and in the reversed direction as, $I_{c,-}\equiv\text{max}_{q>0}\{I(q,\theta+\pi)\}$. The diode efficiency is then defined as,
@@
Here, we defined $\eta$ as non-negative. However, we remark that the modified efficiency $\eta_{\text{signed}}=\frac{I^{+}_{c}-I^{-}_{c}}{I^{+}_{c}+I^{-}_{c}}$ can exhibit sign changes when transitioning through $\eta=0$ [58].
@@
We begin by evaluating the angular dependence of the critical currents and diode efficiencies for fixed $\Delta_{1}$ and $n_{e}$. Our results are shown in Fig. 3(c) and (d) for a Fermi surface with annulus topology. We find that, in general, the diode efficiency takes on finite values. This verifies the emergence of an intrinsic SDE from a fully self-consistent calculation. Furthermore, we find that $\eta$ exhibits a characteristic multi-lobe structure. Specifically, we see that the efficiency tends to be maximized along the directions $\theta\in\{0^{\circ},60^{\circ},120^{\circ}\}$, reaching values of $\sim 30\%$ for the annulus Fermi surface. In addition, it is strongly suppressed along the directions $\theta\in\{30^{\circ},90^{\circ},150^{\circ}\}$. Moreover, the efficiency in the $\theta=0^{\circ},90^{\circ}$ lobes is functionally different from the remaining lobes. We attribute this difference to the small but finite Cooper-pair momentum along $q_{x}$.
@@
It is now interesting to compare these results for the annulus Fermi surface with other Fermi surface topologies. Our results for a single-pocket Fermi surface and a four-pocket Fermi surface are shown in Fig. 4(a) and (b). We see that the shape of the lobes changes qualitatively. Specifically, the width of the lobes around $\theta\in\{60^{\circ},120^{\circ}\}$ is significantly reduced. Furthermore, in these lobes, the maximum efficiency is achieved slightly away from $\theta=60^{\circ},120^{\circ}$. The overall highest efficiency is found for the four-pocket Fermi surface at $\sim 60\%$. These results suggest that the Fermi surface topology can affect the angular dependence of the diode efficiency and, as a result, could provide insights into the type of Fermi surface topology from which the superconducting state arises. We provide more detailed plots of the free energies and critical currents for single-pocket and four-pocket Fermi surface in the Supplemental Material [58].
@@
Finally, we have also investigated the diode efficiency as a function of the diplacement field, $\Delta_{1}$, for different densities, $n_{e}$, and along the fixed $q_{x}$-direction. Our results are shown in Fig. 4(c). We find that, in general, that the diode efficiency seems to increase with increasing displacement field. However, we remark that the system can enter a metallic phase if the displacement field is increased beyond a density-dependent threshold.
Conclusion. In this work, we have theoretically demonstrated and studied the emergence of an intrinsic SDE and non-reciprocal superconductivity in RTLG through a fully self-consistent approach to determining the superconducting order parameter. Our results suggest that angular dependence of the diodicity in RTLG could provide insights into the type of Fermi surface topology from which superconductivity arises.
@@
Acknowledgement. We acknowledge helpful discussions with Marco Valentini.
@@
Yinqi Chen, Constantin Schrade1
@@ 1 Angular-dependence of the diode efficiency
In this first section of the Supplementary Material, we provide more details on the angular dependence of the diode efficiency for different Fermi surfaces. We also suggest a possible experimental setup to probe the angular dependence of the diodicity.
@@ 2 Mean-field decoupling
In this second section of the Supplementary Material, we provide a detailed discussion of the mean-field decoupling of the effective interaction Hamiltonian.
@@ 3 Bogoliubov-de Gennes Hamiltonian
In this third section of the Supplemental Material, we provide details on the derivation of the Bogoliubov-de Gennes Hamiltonian for rhombohedral tetralayer graphene and its diagonalization procedure.
@@ 4 Self-consistency Equation for the Superconducting Gap Function
In this section of the Supplemental Material, we provide details on the derivation of the self-consistency equation that determines the superconducting gap function.
@@ 5 Total Electron Density
In this section of the Supplemental Material, we provide details on the derivation of the total electron density in the superconducting state of the rhombohedral tetralayer graphene system.

### v2

@@ Abstract
One experimental signature of the asymmetric quasiparticle dispersion is that a transparent normal-metal–superconductor junction can exhibit a current-voltage ($I$-$V$) characteristic that is not invariant under $V\rightarrow-V$ [57]. Our results suggest the following probe of the nonreciprocal superconductivity in RTLG: connect two normal leads to the RTLG superconductor, one along $\theta=0^{\circ}$ ($\Gamma$-$K$ direction) and one along $\theta=90^{\circ}$ ($\Gamma$-$M$ direction). Our results imply that the lead along $\theta=0^{\circ}$ would show an asymmytric $I$-$V$ curve, while the $I$-$V$ curve along $\theta=90^{\circ}$ would remain symmetric.
@@
Intrinsic SDE. Next, we address the intrinsic SDE. For this purpose, we self-consistently solve for the superconducting order parameter for a broad range of Cooper-pair momenta, $\boldsymbol{q}$, and use the obtained $\Delta_{\boldsymbol{k},\boldsymbol{q}}$ to evaluate the free energy. The results of the free energy are shown in Fig. 3(a). The current in the ground state that minimizes $F(\mathbf{q})$ is always zero. However, applying a finite current-bias, produces a state with a finite Cooper-pair momentum different from the ground state value, $\boldsymbol{q}\neq\boldsymbol{q}_{0}$. The current in this state is $I=2e|\partial_{\boldsymbol{q}}F(\boldsymbol{q})|$[62], see Fig. 3(b).
@@
We then fix a direction $\theta$, so that $\boldsymbol{q}=(q\cos\theta,q\sin\theta)$ and determine the critical currents in the forward direction, $I^{+}_{c}(\theta)=I_{c}(\theta)$ and in the reversed direction, $I^{-}_{c}(\theta)=I_{c}(\theta+\pi)$. The diode efficiency is defined as,
@@
Here, $\eta$ is non-negative. However, we remark that the difference in critical currents, $\Delta I_{c}=I^{+}_{c}-I^{-}_{c}$, can exhibit sign changes when transitioning through $\eta=0$, see [63].
@@
We begin by evaluating the angular dependence of the critical currents and diode efficiencies for fixed $\Delta_{1}$ and $n_{e}$. Our results are shown in Fig. 3(c) and (d) for a Fermi surface with annulus topology. We find that, in general, the diode efficiency takes on finite values. This verifies the emergence of an intrinsic SDE from a fully self-consistent calculation. Furthermore, we find that $\eta$ exhibits a characteristic multi-lobe structure. Specifically, we see that the efficiency tends to be maximized along the directions $\theta\in\{0^{\circ},60^{\circ},120^{\circ}\}$, reaching values of $\sim 30\%$ for the annulus Fermi surface. In addition, as a result of the reflection symmetry $\sigma_{v}$, $\eta$ vanishes along the directions perpendicular to one of the three mirror planes [64], i.e., for $\theta\in\{30^{\circ},90^{\circ},150^{\circ}\}$.
@@
It is interesting to compare these results to other Fermi surface topologies. In Fig. 4(a–d), we present our findings for both a single-pocket and a four-pocket Fermi surface. As shown in Fig. 4(a) and (b), the angular dependence of the forward and reverse critical currents, $I_{c}^{\pm}$, differs markedly between the two cases. The single-pocket Fermi surface exhibits sharp peaks in $I_{c}^{\pm}$ that are absent for the four-pocket topology. These differences are also reflected in the diode efficiencies shown in Fig. 4(c) and (d), which are more sharply peaked in the single-pocket case. Compared to the previous annular Fermi surface discussed in Fig. 3, the lobes for both the single-pocket and four-pocket cases also appear narrower. Altogether, these results indicate that the Fermi surface topology strongly influences the angular dependence of both the critical currents and the diode efficiency. In particular, the angular profile of the diode efficiency and the critical currents could provide insights into the type of Fermi surface topology from which the superconducting state arises. More detailed plots of the free energies and supercurrents for single- and four-pocket Fermi surfaces are shown in the Supplemental Material [63].
@@
Finally, we also briefly comment on the dependence of the diode efficiency on $\Delta_{1}$. Our results of the diode efficiencies along the $x$-direction versus $\Delta_{1}$ for different densities, $n_{e}$, are shown in Fig. 4(c). We find that the diode efficiency tends to increase with increasing $\Delta_{1}$. However, we also observe that when increasing $\Delta_{1}$ above a threshold, the superconducting gap closes and the system enters a metallic phase.
Conclusion. We have theoretically studied the intrinsic SDE as a novel probe of the candidate chiral superconducting state in RTLG. In particular, we have argued that the angular dependence of the critical current asymmetry and of the SDE in this system can provide a route to determine the Fermi surface topology from which superconductivity arises. We note that these angular dependencies can be probe experimentally using sunflower-shaped samples [30, 65]. We hope that our work will contribute to understanding novel superconducting states in RTLG and to the development of novel techniques for their detection [66].
@@
Acknowledgement. We acknowledge helpful discussions with Marco Valentini and Sayan Banerjee.
@@
Yinqi Chen1, Mathias S. Scheurer2, Constantin Schrade1
@@
2Institute for Theoretical Physics III, University of Stuttgart, 70550 Stuttgart, Germany
@@ 1 Angular-dependence of the diode efficiency
Here, we provide details on the angular dependence of the diode efficiency for different Fermi surfaces. We also suggest a possible experimental setup to probe the angular dependence of the diodicity.
@@ 2 Mean-field decoupling
Here, we provide a detailed discussion of the mean-field decoupling of the effective interaction Hamiltonian.
@@ 3 Bogoliubov-de Gennes Hamiltonian
Here, we provide details on the derivation of the Bogoliubov-de Gennes Hamiltonian for rhombohedral tetralayer graphene and its diagonalization procedure.
@@ 4 Self-consistency Equation for the Superconducting Gap Function
Here, we provide details on the derivation of the self-consistency equation that determines the superconducting gap function.
@@ 5 Total Electron Density
Here, we provide details on the derivation of the total electron density in the superconducting state of the rhombohedral tetralayer graphene system.